OpenStudy (ms2016):

A 22 gram sample of a substance that's used to sterilize surgical instruments has a k-value of 0.138. N=N0e^-kt N0=Initial mass (at time t=0) N=mass at time t k= a positive constant that depends on the substance itself and on the units used to measure time t=time, in days What is the substance's half-life, in days rounded to the nearest tenth?

OpenStudy (holsteremission):

The half-life is the time it takes for the substance to decay to half its original amount. This means if the half-life is \(T\), then \(N=\dfrac{N_0}{2}\) and the half-life is given by \[\frac{1}{2}=e^{-kT}\implies T=\frac{1}{k}\ln2\]Plug in \(k\) and you get your answer.